Check my work Calculate the present value of the following annuity streams: a. $5,000 received each year for 6 years on the last day of each year if your investments pay 6 percent compounded annually b. $5,000 received each quarter for 6 years on the last day of each quarter if your investments pay 6 percent compounded quarterly c. $5,000 received each year for 6 years on the first day of each year if your investments pay 6 percent compounded annually d. $5,000 received each quarter for 6 years on the first day of each quarter if your investments pay 6 percent compounded quarterly. For all requirements, do not round intermediate calculations. Round your answers to 2 decimal places. (e.g., 32.16)) a. Present value b. Present value c. Present value d. Present value
Answer #1

a) PV of annuity = 5000*(1-(1+6%)-6)/6% = 24,586.62

b) PV of annuity rate per quarter = 5000*(1-(1+6%/4)-24/(6%/4) = 100,152.03

c) PV of annuity due = (1+6%)*(5000*(1-(1+6%)-6)/(6%)= 26,061.82

d) PV of annuity due rate per quarter = (1+6%/4)*(5000*(1-(1+6%/4)-24/(6%/4) = 101,654.31

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well